OpenStudy (anonymous):

log x^2-2x-3/x^2-4-log x^2+7x+6/x+2 epress as single log

OpenStudy (anonymous):

x^2 -2x -3 = (x-3)(x+1) x^2-4 =(x-2)(x+2) x^2 +7x +6 = (x+6)(x+1)

OpenStudy (saifoo.khan):

.

OpenStudy (anonymous):

\[\log (A) -\log(B) = \log \frac{A}{B}\]

OpenStudy (anonymous):

\[\frac{A}{B} = \frac{ (x-3)(x+1)}{(x+2)(x-2)} \times \frac{x+2}{(x+6)(x+1)}\]

OpenStudy (anonymous):

\[\frac{A}{B} = \frac{x-3}{(x-2)(x+6)}\]

OpenStudy (anonymous):

Thanks could YOU help me get some more

OpenStudy (anonymous):

\[\log \frac{(x-3)}{(x-2)(x+6)}\]